# TreeRare: Syntax Tree-Guided Retrieval and Reasoning for Knowledge-Intensive Question Answering

### v1

@@ Abstract
In real practice, questions are typically complex and knowledge-intensive, requiring Large Language Models (LLMs) to recognize the multifaceted nature of the question and reason across multiple information sources. Iterative and adaptive retrieval, where LLMs decide when and what to retrieve based on their reasoning, has been shown to be a promising approach to resolve complex, knowledge-intensive questions. However, the performance of such retrieval frameworks is limited by the accumulation of reasoning errors and misaligned retrieval results. To overcome these limitations, we propose TreeRare (Syntax Tree-Guided Retrieval and Reasoning), a framework that utilizes syntax trees to guide information retrieval and reasoning for question answering. Following the principle of compositionality, TreeRare traverses the syntax tree in a bottom-up fashion, and in each node, it generates subcomponent-based queries and retrieves relevant passages to resolve localized uncertainty. A subcomponent question answering module then synthesizes these passages into concise, context-aware evidence. Finally, TreeRare aggregates the evidence across the tree to form a final answer. Experiments across five question answering datasets involving ambiguous or multi-hop reasoning demonstrate that TreeRare achieves substantial improvements over existing state-of-the-art methods.
@@ 4.1 Experiment Setup
Baselines.We evaluate TreeRare against a comprehensive suite of baselines that represent key prompting and planning paradigms under a unified retriever backbone (BM25 Robertson et al. (2009)). (1) Zero-shot and few-shot prompting is introduced by Brown et al. (2020). These serve as foundational setups without any intermediate reasoning steps. (2) Chain-of-Thought prompting (CoT) Wei et al. (2022) encourages step-by-step reasoning by appending an instruction such as “Let’s think step by step” to the input. (3) Self-Consistency (SC) Wang et al. (2022) samples multiple reasoning paths and selects the final answer via majority voting. (4) Tree-of-Thoughts (ToT) Yao et al. (2023) explores multiple structured reasoning trajectories using tree-based search and pruning strategies. (5) ReAct Yao et al. (2022) interleaves reasoning and retrieval by prompting the model to decide dynamically when and what to retrieve. We adapt ReAct to use BM25 instead of web-based tools, denoted as ReAct*. Details of the baseline implementations are included in Appendix B.
@@
Table 1: Performance of TreeRare (CT) TreeRare(DT), Chain-of-Thought (CoT), Self-Consistency (CoT-SC), ReAct, ReAct*, and Tree-of-Thoughts (ToT) across five different QA datasets. AVG indicates the average COV-EM on the three multihop datasets. ReAct* denotes a BM25-based variant of ReAct. Bold marks the best performance, and underline denotes the second-best under teh same setting. Rel. Impr. stands for relative improvement over the best baseline in the same setting.
@@ 4.2 Results
COV-EM COV-EM COV-EM AR ER COV-EM Dis-F1
BM25 0.473 0.151 0.520 0.381 0.409 0.539 0.423 0.328
DPR 0.392 0.132 0.318 0.281 0.343 0.449 0.438 0.343
BM25+DPR 0.456 0.146 0.380 0.327 0.322 0.404 0.480 0.351
Tree-Retrieval (DT) 0.528 0.138 0.562 0.409 0.427 0.581 0.514 0.361
Tree-Retrieval (CT) 0.510 0.156 0.538 0.401 0.558 0.681 0.533 0.335
@@ 5.1 Tree-Retrieval
As shown in Table 2, Tree-Retrieval consistently surpasses classical retrieval approaches such as BM25, DPRKarpukhin et al. (2020), and their hybrid combination (BM25+DPR) across all QA datasets. The performance gain shows that incorporating syntactic structure into the retrieval process enhances the relevance of retrieved documents. Additionally, these results suggest that the effectiveness of TreeRare cannot be solely attributed to enhanced downstream reasoning by the LLMs. Rather, a significant portion of its advantage stems from its retrieval stage, which is structurally guided to extract more fine-grained, contextually aligned documents.
@@ 5.3 Error Analysis
To better understand the performance differences between ReAct and TreeRare, we conduct a human evaluation comparing their outputs on randomly sampled multihop questions. Each output is manually categorized into one of five distinct outcome types: Success, Retrieval Error, Reasoning Error, Partial Answer, and Label Ambiguity. The detailed definitions and illustrative examples of each error type are presented in Appendix F.
@@ Appendix A TreeRare Implementation Details
Traversing Details In the experiment, TreeRare conducts a pre-order traversal. Each node is processed only after all its child nodes are resolved. We maintain a processing queue initialized with leaf nodes. When processing a node: If it has no children, we directly use its text span to generate quries. If it has children, we first aggregate evidence from its children before proceeding to subcomponent-specific query generation. If it has no children, we directly use its text span as the initial evidence. To improve computational efficiency, we implement a pruning mechanism that skips nodes based on two criteria: Nodes with syntactic types typically considered non-informative (e.g., punctuation, determiners, conjunctions). Nodes whose associated sub-phrases are shorter than three tokens.
Subcomponent Query Generation. At each non-leaf node, we generate queries to resolve syntactic uncertainty associated with its sub-phrase. We prompt the LLM using templates detailed in Appendix G We generate up to five candidate queries and select the top three queries according to heuristic rules prioritizing coverage and specificity.
@@ Appendix E Further Cost Analysis
We measure the API cost of GPT-4o-mini by summing the number of input and output tokens processed for each method across five datasets: HotpotQA, MuSiQue, 2WikiMultihopQA, AmbigDoc, and ASQA. For each dataset, 500 examples were randomly sampled and processed with each method. The input and output tokens were multiplied by OpenAI’s published pricing for GPT-4o-mini (as of April 2025).
Appendix F Case Study
@@
we randomly sampled 50 multihop questions from the HotpotQA development set. For each sample, a human annotator examined the reasoning trace produced by both ReAct and TreeRare, along with the final answer and supporting documents. We present one representative example for each failure category.
@@ Appendix G Prompt for TreeRare
Final answer generation prompt for multi hop QA Answer the following question: {{questions}} , with following documents: {{documents}}. Your response should strictly follow the format: Explanations :[give your step by step Analysis here ] FINAL:(BE CONCISE, ONLY a FEW phrases) let’s think step by step
@@
Subcomponent Question generation prompt for Ambiguous QA You’re a disambiguation expert analyzing "{{phrase}}" in: {{self.question}} Instruction: 1. Analyze the question by considering these potential ambiguities: - Temporal: Check for unclear time references, periods, or temporal scope - Entity: Identify names, references, or terms that could refer to multiple entities - Semantic: Look for words with multiple meanings (polysemy/homonymy) - Scope: Consider possible boundaries and levels of detail - Intent: Examine possible purposes and expected answer types - Cultural: Consider cultural-dependent interpretations - Quantitative: Check for unclear measurements or numerical references - Linguistic: Analyze syntax and referential clarity - Categorical: Consider possible classification schemes - Contextual: Examine required background knowledge and relationships 2. Analyze the question word by word. Return disambiguated question and its interperatation for each different meaning Here is what we currently know Documents:{{context}} pick top 5 questions that are best in disambiguating the question. (covers different meanings of the questions) and strictly FOLLOW the format: response: question1; question2;….
@@
Final answer generation prompt for Ambiguous QA The question may be ambiguous and have multiple correct answers, and in that case, you have to provide a long-form answer including all correct answers. 1. Carefully go through all the given documents. 2.The using your and context, provide answer. Your response should strictly follow the format: Explanations (Step 2):[give your step by step Analysis here ] FINAL(Step 2): Please ONLY reply according to this format Question: {{questions}} Document: {{documents}} let’s think step by step
@@
Subcomponent question answering prompt Answer the {{questions}} based on on the document info. For each question find as many answers as possible. Response all the answers in a short paragraph (as specific as possible). Relevant Document: {{context}}
Subcomponent Question generation prompt for multihop QA You’re a an expert analyzing "{{phrase}}" in main query: "{{self.question}}". 1.You should generate at most 5 simple questions that mainly ask for information about {{phrase}} that helps understand the main query. Question should be single-hop, clear and search-friendly. 2. Here is what we currently know Documents:{{context}} pick top 3 questions that are best in decomposing the main query and leading to final answer of the main query. strictly FOLLOW the format: response: question1; question2;….

### v2

@@ Abstract
In real practice, questions are typically complex and knowledge-intensive, requiring Large Language Models (LLMs) to recognize the multifaceted nature of the question and reason across multiple information sources. Iterative and adaptive retrieval, where LLMs decide when and what to retrieve based on their reasoning, has been shown to be a promising approach to resolve complex, knowledge-intensive questions. However, the performance of such retrieval frameworks is limited by the accumulation of reasoning errors and misaligned retrieval results. To overcome these limitations, we propose TreeRare (Syntax Tree-Guided Retrieval and Reasoning), a framework that utilizes syntax trees to guide information retrieval and reasoning for question answering. Following the principle of compositionality, TreeRare traverses the syntax tree in a bottom-up fashion, and in each node, it generates subcomponent-based queries and retrieves relevant passages to resolve localized uncertainty. A subcomponent question answering module then synthesizes these passages into concise, context-aware evidence. Finally, TreeRare aggregates the evidence across the tree to form a final answer. Experiments across five question answering datasets involving ambiguous or multi-hop reasoning demonstrate that TreeRare achieves substantial improvements over existing state-of-the-art methods. [^1] Our code is publicly available at https://github.com/billycrapediem/TreeRare.
@@ 4.1 Experiment Setup
Baselines.We evaluate TreeRare against a comprehensive suite of baselines that represent key prompting and planning paradigms under a unified retriever backbone (BM25 Robertson et al. (2009)). (1) Zero-shot and few-shot prompting is introduced by Brown et al. (2020). These serve as foundational setups without any intermediate reasoning steps. (2) Chain-of-Thought prompting (CoT) Wei et al. (2022) encourages step-by-step reasoning by appending an instruction such as “Let’s think step by step” to the input. (3) Self-Consistency (SC) Wang et al. (2022) samples multiple reasoning paths and selects the final answer via majority voting. (4) Tree-of-Thoughts (ToT) Yao et al. (2023) explores multiple structured reasoning trajectories using tree-based search and pruning strategies. (5) ReAct Yao et al. (2022) interleaves reasoning and retrieval by prompting the model to decide dynamically when and what to retrieve. We adapt ReAct to use BM25 instead of web-based tools, denoted as ReAct*. (6) Topology-of-Question-Decomposition (ToQD) Li et al. (2025) constructs a topology graph of sub-questions and uses self-verify inference to selectively activate retrieval only when necessary. Details of baselines implementations are included in Appendix B.
@@
Table 1: Performance of TreeRare (CT) TreeRare(DT), Chain-of-Thought (CoT), Self-Consistency (CoT-SC), ReAct, ReAct*, Tree-of-Thoughts (ToT) and Topology-of-Question-Decomposition (ToQD) across five different QA datasets. AVG indicates the average COV-EM on the three multihop datasets. ReAct* denotes a BM25-based variant of ReAct. Bold marks the best performance, and underline denotes the second-best under teh same setting. Rel. Impr. stands for relative improvement over the best baseline in the same setting.
@@ 4.2 Results
COV-EM COV-EM COV-EM AR ER Dis-F1 COV-EM
BM25 0.473 0.151 0.520 0.381 0.409 0.539 0.328 0.423
DPR 0.392 0.132 0.318 0.281 0.343 0.449 0.343 0.438
BM25+DPR 0.456 0.146 0.380 0.327 0.322 0.404 0.351 0.480
IRCoT 0.486 0.148 0.504 0.379 0.314 0.394 0.335 0.488
Tree-Retrieval (DT) 0.528 0.138 0.562 0.409 0.427 0.581 0.361 0.514
Tree-Retrieval (CT) 0.510 0.156 0.538 0.401 0.558 0.681 0.335 0.533
@@ 5.1 Tree-Retrieval
As shown in Table 2, Tree-Retrieval consistently surpasses classical retrieval approaches such as BM25, DPRKarpukhin et al. (2020), and LLM based question decomposition approaches like IRCoT Trivedi et al. (2023) across all QA datasets. The performance gain shows that incorporating syntactic structure into the retrieval process enhances the relevance of retrieved documents. Additionally, these results suggest that the effectiveness of TreeRare cannot be solely attributed to enhanced downstream reasoning by the LLMs. Rather, a significant portion of its advantage stems from its retrieval stage, which is structurally guided to extract more fine-grained, contextually aligned documents.
@@ 5.3 Error Analysis
To better understand the performance differences between ReAct and TreeRare, we conduct a human evaluation comparing their outputs on randomly sampled multihop questions. Each output is manually categorized into one of five distinct outcome types: Success, Retrieval Error, Reasoning Error, Partial Answer, and Label Ambiguity. The detailed definitions and illustrative examples of each error type are presented in Appendix G.
@@ Appendix A TreeRare Implementation Details
Traversing Details In the experiment, TreeRare conducts a pre-order traversal. Each node is processed only after all its child nodes are resolved. We maintain a processing queue initialized with leaf nodes. When processing a node: If it has no children, we directly use its text span to generate quries. If it has children, we first aggregate evidence from its children before proceeding to subcomponent-specific query generation. If it has no children, we directly use its text span as the initial evidence. To improve computational efficiency, we implement a pruning mechanism that skips nodes based on two criteria: Nodes with syntactic types typically considered non-informative (e.g., punctuation, determiners, conjunctions). Nodes whose associated sub-phrases are minimum phrase length ($L_{min}$). We sets $L_{min}=3$ for all the experiments in the paper. In
Subcomponent Query Generation. At each non-leaf node, we generate queries to resolve syntactic uncertainty associated with its sub-phrase. We prompt the LLM using templates detailed in Figure10, Figure12, Figure13, Figure14. We generate up to five candidate queries and select the top three queries according to heuristic rules prioritizing coverage and specificity.
@@ Appendix B Baseline Implementation Details
Topology-of-Question-Decomposition. We follow the implementation of the ToQD paper. To make sure every method uses the same retriever, we modify the retrieval module of ToQD into the same as the retrieval module in TreeRare.
@@ Appendix C Ablation Study Details
$L_{min}$ HotpotQA MuSiQue 2WikiMQA AVG AmbigDoc ASQA
COV-EM COV-EM COV-EM AR ER COV-EM Dis-F1
3 0.544 0.240 0.583 0.457 0.592 0.722 0.381 0.547
6 0.531 0.224 0.547 0.434 0.542 0.703 0.333 0.441
10 0.512 0.198 0.492 0.318 0.504 0.600 0.328 0.448
Table 5: Impact of minimum phrase length ($L_{min}$) on TreeRare (DT) performance across across HotpotQA, MuSiQue, 2WikiMQA, AmbigDoc, and ASQA based on OpenAI pricing.
@@ Appendix E Further Cost Analysis
As shown in Figure5, we measure the API cost of GPT-4o-mini by summing the number of input and output tokens processed for each method across five datasets: HotpotQA, MuSiQue, 2WikiMultihopQA, AmbigDoc, and ASQA. For each dataset, 500 examples were randomly sampled and processed with each method. The input and output tokens were multiplied by OpenAI’s published pricing for GPT-4o-mini (as of April 2025).
Appendix F Further Analysis of Minimum Phrase Length
To analysis the impact of minimum phrase length of TreeRare, we performs additional experiment by setting $L_{min}$ to $6$ and $10$. As shown in 5, TableTreeRare demonstrates consistently superior performance with $L_{min}=3$ across most evaluation metrics. The performance degradation with increased $L_{min}$ values suggests that finer-grained syntactic analysis enables more precise identification of knowledge gaps and uncertainty points within complex questions. Figure 6 reveals the computational trade-offs associated with different $L_{min}$. $L_{min}$ values result in higher API costs due to increased node processing. This cost increase stems from processing more nodes in the syntax tree, each requiring query generation, retrieval, and subcomponent question answering.
Figure 6: Total GPT-4o-mini API cost (input + output tokens) for TreeRare(CT) on different minimum phrase lengths (3,6,10) across HotpotQA, MuSiQue, 2WikiMQA, AmbigDoc, and ASQA based on OpenAI pricing.
Appendix G Case Study
@@
we randomly sampled 50 multihop questions from the HotpotQA development set. For each sample, a human annotator examined the reasoning trace produced by both ReAct and TreeRare, along with the final answer and supporting documents. We present one representative example for each failure category. Additionally, we provide a single case study of concrete output TreeRare on one single question at Figure15 and Figure16.
@@
Answer the following question: {{questions}} , with following documents: {{documents}}. Your response should strictly follow the format: Explanations :[give your step by step Analysis here ] FINAL:(BE CONCISE, ONLY a FEW phrases) let’s think step by step
Figure 11: Final answer generation prompt for multi hop QA
@@
You’re a disambiguation expert analyzing "{{phrase}}" in: {{self.question}} Instruction: 1. Analyze the question by considering these potential ambiguities: - Temporal: Check for unclear time references, periods, or temporal scope - Entity: Identify names, references, or terms that could refer to multiple entities - Semantic: Look for words with multiple meanings (polysemy/homonymy) - Scope: Consider possible boundaries and levels of detail - Intent: Examine possible purposes and expected answer types - Cultural: Consider cultural-dependent interpretations - Quantitative: Check for unclear measurements or numerical references - Linguistic: Analyze syntax and referential clarity - Categorical: Consider possible classification schemes - Contextual: Examine required background knowledge and relationships 2. Analyze the question word by word. Return disambiguated question and its interperatation for each different meaning Here is what we currently know Documents:{{context}} pick top 5 questions that are best in disambiguating the question. (covers different meanings of the questions) and strictly FOLLOW the format: response: question1; question2;….
Figure 12: Subcomponent Question generation prompt for Ambiguous QA
@@
The question may be ambiguous and have multiple correct answers, and in that case, you have to provide a long-form answer including all correct answers. 1. Carefully go through all the given documents. 2.The using your and context, provide answer. Your response should strictly follow the format: Explanations (Step 2):[give your step by step Analysis here ] FINAL(Step 2): Please ONLY reply according to this format Question: {{questions}} Document: {{documents}} let’s think step by step
Figure 13: Final answer generation prompt for Ambiguous QA
@@
Answer the {{questions}} based on on the document info. For each question find as many answers as possible. Response all the answers in a short paragraph (as specific as possible). Relevant Document: {{context}}
Figure 14: Subcomponent question answering prompt
example Output of TreeRare Question: What screenwriter with credits for ‘Evolution’ co-wrote a film starring Nicolas Cage and Téa Leoni? Dependency Parser Output:
ROOT []
|   co-wrote --root--> [What screenwriter with credits for
"Evolution" co-wrote a film starring
Nicolas Cage and Tea Leoni]
|  |-- screenwriter --nsubj--> [screenwriter with credits for "Evolution"]
|  |  |  |-- What --det--> [What]
|  |  |  -- credits --nmod--> [credits for "Evolution"]
|  |  |     |  |-- with --case--> [with]
|  |  |     |  -- Evolution --nmod--> [Evolution]
|  |-- film --obj--> [film starring Nicolas Cage and Tea Leoni]
|  |  |  |-- a --det--> [a]
|  |  |  -- starring --acl--> [starring Nicolas Cage and Tea Leoni]
|  |  |     |  -- Nicolas Cage --obj--> [Nicolas Cage and Tea Leoni]
|  |  |     |     |  -- Tea Leoni --conj--> [Tea Leoni]
|  -- ? --punct--> [?]
Figure 15: Output of TreeRare at question decomposition stage
example Output of TreeRare Node: credits for ‘Evolution’ ” Generated Questions: Who wrote the screenplay for the film Evolution”? What movie starring Nicolas Cage and Tea Leoni was co-written by the same screenwriter as Evolution”? What year was Evolution” released? Subcomponent Question Answering: The screenplay for the film Evolution” was co-written by David Diamond and David Weissman, along with a story by Don Jakoby. The same screenwriters, Diamond and Weissman, also co-wrote the film The Family Man,” which stars Nicolas Cage and Tea Leoni. “Evolution” was released in the year 2001. Node: screenwriter with credits for ‘Evolution’ ” Generated Questions: Who are the screenwriters credited for Evolution”? What is the title of the film starring Nicolas Cage and Tea Leoni co-written by the same screenwriters? In what year was “Evolution” released? Subcomponent Question Answering: The screenwriters credited for Evolution” are David Diamond and David Weissman. They also co-wrote the film The Family Man,” which stars Nicolas Cage and Téa Leoni. “Evolution” was released in the year 2001. Node: Nicolas Cage and Téa Leoni” Generated Questions: What movies have Nicolas Cage and Téa Leoni starred in together? Who co-wrote the screenplay for the film Evolution”? What are the main themes of the film “Evolution”? Subcomponent Question Answering: Nicolas Cage and Téa Leoni starred together in the film The Family Man,” a romantic comedy-drama released in 2000. The screenplay for Evolution” was co-written by David Diamond and David Weissman, in collaboration with Don Jakoby, who originally came up with the story. The main themes of “Evolution” include the unintended consequences of technological advancement and the humorous exploration of evolution itself, particularly focusing on alien organisms and how they adapt to Earth, leading to unexpected and comedic situations. Node: starring Nicolas Cage and Téa Leoni” Generated Questions: What film did Nicolas Cage and Téa Leoni star in together? Who were the co-writers of the screenplay for Evolution”? What are the main themes explored in the film Evolution”? When was The Family Man” released? What genre does the film “The Family Man” belong to? Subcomponent Question Answering: Nicolas Cage and Téa Leoni starred together in the film The Family Man,” which was released in 2000. The screenplay for Evolution” was co-written by David Diamond and David Weissman. The film Evolution” explores themes such as the rapid development and adaptation of species, the impact of science and technology on society, and environmental issues tied to evolution. The Family Man” was released on December 22, 2000, and it belongs to the romantic comedy-drama genre. Node: film starring Nicolas Cage and Téa Leoni” Generated Questions:What film did Nicolas Cage and Téa Leoni star in together? Who were the co-writers of the screenplay for Evolution”?When was “The Family Man” released? Subcomponent Question Answering: Nicolas Cage and Téa Leoni starred together in the film The Family Man,” which was released in 2000. The screenplay for Evolution” was co-written by David Diamond and David Weissman, who frequently collaborate on various projects. “The Family Man” premiered on December 22, 2000. Final Answer: The screenwriters credited for the film Evolution” are David Diamond and David Weissman.
Figure 16: Output of TreeRare at the Subcomponent Question Answering Stage